5. The NPV and payback period

What information does the payback period provide?

Suppose you are evaluating a project with the expected future cash inflows shown in the following table. Your boss has asked you to calculate the project’s net present value (NPV). You don’t know the project’s initial cost, but you do know the project’s regular, or conventional, payback period is 2.50 years.

Year

Cash Flow

Year 1 $350,000
Year 2 $475,000
Year 3 $400,000
Year 4 $475,000

If the project’s weighted average cost of capital (WACC) is 9%, the project’s NPV (rounded to the nearest dollar) is:

$273,019

$392,465

$341,274

$358,338

Which of the following statements indicate a disadvantage of using the regular payback period (not the discounted payback period) for capital budgeting decisions? Check all that apply.

The payback period is calculated using net income instead of cash flows.

The payback period does not take the time value of money into account.

The payback period does not take the project’s entire life into account.
Answer #1

Payback period=Full years until recovery+(Unrecovered cost at the beginning of last year)/(Cash flow during the last year)

Given that the payback period is 2.5, it means it took 2.5 years for the initial amount to get recivered
So, the amount will be the cash inflows in first 2 years and half of the amount in year 3
Initial cash flow amount=350000+475000+(0.5)*400000=1025000

NPV=-Initial cash outflow + Present value of cash inflows
=-1025000 + 350000/(1+9%)^1+475000/(1+9%)^2+400000/(1+9%)^3+475000/(1+9%)^4
=-1025000 + 321100.9174+399797.9968+308873.392+336501.9753
=341274.28 or 341274 (Rounded to the nearest whole number)

So, the NPV should be $341274

The two disadvantages of simple payback period method are:
The payback period does not take the time value of money into account.
The payback period does not take the project’s entire life into account.